OpenStudy (unklerhaukus):

\[(x-2y+1)\text dx+(4x-3y-6)\text dy=0\] \[\frac{\text dy}{\text dx}=\frac{2y-x-1}{4x-3y-6}\]

OpenStudy (unklerhaukus):

This

OpenStudy (unklerhaukus):

i have clearly mad aan error

OpenStudy (unklerhaukus):

error(s)

OpenStudy (sasogeek):

differentiate implicitly? are you asking for it to be checked for you or... i'm not sure what the question is...

OpenStudy (unklerhaukus):

there is mistakes in my working because wolfram 's solution is different to mine, can you what i have done wrong

OpenStudy (sasogeek):

what was your solution?

OpenStudy (unklerhaukus):

it is in pdf 9.pdf, up the page

OpenStudy (sasogeek):

ohhh i see, so the question is what you put in the wolfram link and your answer is what's in the pdf link... right?

OpenStudy (unklerhaukus):

yes, the question is also in the blue question box at the top of the page

OpenStudy (sasogeek):

i'm not sure what we're supposed to do here... are we supposed to solve to the second derivative?

OpenStudy (sasogeek):

or we're supposed to integrate?

OpenStudy (unklerhaukus):

there is not second derivative, the question has a form of the first derivative of y with respect to x, solve the differential equation. ie find an expression for y(x) that does not include derivatives

OpenStudy (sasogeek):

sorry but I can't do this integration, i'll see if I can get you some help with @ash2326

OpenStudy (ash2326):

This is a differential equation, I've to check it

OpenStudy (sasogeek):

well i thought anti-derivatives were the same as integration... right? from dy/dx to y(x) = integration and from y(x) to dy/dx=differentiation.... ? :|

OpenStudy (ash2326):

Yeah you're right but we have to make some modifications before we can integrate to find y. @UnkleRhaukus jusr wait for some time

OpenStudy (unklerhaukus):

ok ill wait

OpenStudy (ash2326):

@UnkleRhaukus Sorry I forgot this form. @lalaly might be able to help you

OpenStudy (anonymous):

@UnkleRhaukus, You made a very silly mistake. You wrote dX/dY instead of dY/dX While substituting x=X+h y=Y+k

OpenStudy (unklerhaukus):

i did too, thanks @shivam_bhalla.

OpenStudy (anonymous):

Hope you are getting the answer now :)

OpenStudy (unklerhaukus):

that only changed two terms , that mistake didn't carry through vary far and hasn't changed my result